# Cyclic image generation using chaotic dynamics

### v1

@@ Introduction
MNIST is a widely used dataset consisting of 70000 grayscale images of handwritten digits from 0 to 9. Of these images, 60000 are provided for training and 10000 are provided for testing. Each image has a resolution of $28\times 28$ pixels, resulting in a total of 784 pixels per image. For our experiments with MNIST, we used subsets of images belonging to specific categories: images of the digit 0 were assigned to dataset $X$, images of the digit 1 were assigned to dataset $Y$, and images of the digit 2 were assigned to dataset $Z$ (Fig 1).
Fashion-MNIST [27] is another dataset that serves as a drop-in replacement for MNIST. It consists of 70000 grayscale images from 10 categories of fashion products. The dataset maintains the same image size and split as MNIST, with 60000 training images and 10000 testing images, each having a resolution of $28\times 28$ pixels. We used subsets of images belonging to specific fashion product categories: images of T-shirts/tops were assigned to dataset $X$, images of sneakers were assigned to dataset $Y$, and images of bags were assigned to dataset $Z$ (Fig 1).
@@ Results
To visually assess the distribution of the image sequences generated by the iteration of $G$, we visualize the distribution using UMAP [28]. We iterated the transformation 5000 times from an initial image and visualized the distribution of the 5000 generated data points on a 2D plane (Fig 4 left and right). To embed, we first trained the UMAP model to embed the training data distribution into a 2D latent space, then we embedded the generated data into the same latent space using the learned model. In the figures, the training data distributions of datasets $X$, $Y$, and $Z$ (digits 0, 1, and 2 for MNIST and T-shirt/top, sneaker, and bag for Fashion-MNIST) are represented by cyan, pink, and green points, respectively, while the generated data are represented by purple points.
@@
To quantitatively evaluate the quality and diversity of the generated image sequences, we calculated the precision and recall [19] (P/R) of the generated images from the MNIST-trained model using the method described in the subsection entitled “Precision and recall.” Here, quality refers to how similar the generated images are to the real dataset images, whereas diversity refers to how well the generated images cover a wide range of variations in the real dataset. The important parameters in this method are the number of data samples and the number of nearest neighbors. In this method, an equal number of real and generated data samples should be prepared. In our case, the real data consisted of all the test dataset images from categories $X$, $Y$, and $Z$, amounting to $3178$ samples. For the generated images, transformations were repeatedly applied starting from a test image $x_{0}$ selected from category $X$ as an initial value, resulting in a trajectory of 3178 points $\{x_{1},\ldots,x_{3178}\}$, and these samples included in this trajectory were used for evaluation. Previous research [19] proposed using 20000 samples each for real and generated data as a standard procedure. However, in this case, the number of test images does not reach this number, so direct comparison with other literature results is impossible. To ensure the validity of the evaluation, $3178$ samples from the same categories were extracted from the training dataset, and the P/R between the training and test datasets was calculated as a reference value. In the literature [19], the number of nearest neighbors $k$ is typically set to 3. However, because the number of data samples used in this study was different from that in the literature, it is difficult to draw conclusions from results with only a single value of $k$. Therefore, we varied $k$ from 1 to 10 and observed the resulting changes in P/R values.
@@
S4 Fig. shows the histograms of the first five Lyapunov exponents calculated for each of the 980 trajectories. The histograms exhibit unimodal Gaussian-like distributions, suggesting that these trajectories converge to the same attractor without multi-stability. Therefore, it is reasonable to average these values to estimate the Lyapunov exponents.
Fig 8 shows all the spectrum values, with the inset showing an enlarged view of the first 15 Lyapunov exponents. The first five Lyapunov exponents are definitely positive, and the largest Lyapunov exponent is estimated to be about 0.340. The presence of these positive Lyapunov exponents indicates that the generated trajectories exhibit sensitive dependence on initial conditions and are chaotic, suggesting that the chaotic attractor generates various images.

### v2

@@ Introduction
MNIST is a widely used dataset consisting of $7\times 10^{4}$ grayscale images of handwritten digits from 0 to 9. Of these images, $6\times 10^{4}$ are provided for training and $10^{4}$ are provided for testing. Each image has a resolution of $28\times 28$ pixels, resulting in a total of 784 pixels per image. For our experiments with MNIST, we used subsets of images belonging to specific categories: images of the digit 0 were assigned to dataset $X$, images of the digit 1 were assigned to dataset $Y$, and images of the digit 2 were assigned to dataset $Z$ (Fig 1).
Fashion-MNIST [27] is another dataset that serves as a drop-in replacement for MNIST. It consists of $7\times 10^{4}$ grayscale images from 10 categories of fashion products. The dataset maintains the same image size and split as MNIST, with $6\times 10^{4}$ training images and $10^{4}$ testing images, each having a resolution of $28\times 28$ pixels. We used subsets of images belonging to specific fashion product categories: images of T-shirts/tops were assigned to dataset $X$, images of sneakers were assigned to dataset $Y$, and images of bags were assigned to dataset $Z$ (Fig 1).
@@ Results
To visually assess the distribution of the image sequences generated by the iteration of $G$, we visualize the distribution using UMAP [28]. We iterated the transformation 5000 times from an initial image and visualized the distribution of the 5000 generated data points on a 2D plane (Fig 4 left and right). To embed, we first trained the UMAP model to embed the training data distribution into a 2D latent space, then we embedded the generated data into the same latent space using the learned model. In the figures, the training data distributions of datasets $X$, $Y$, and $Z$ (digits 0, 1, and 2 for MNIST and T-shirt/top, sneaker, and bag for Fashion-MNIST) are represented by cyan, pink, and green points, respectively, while the generated data are represented by purple points. Note that the quantitative analysis in this paper was not performed based on the data in the UMAP-transformed space. The analysis of Lyapunov exponents was performed in the original image space, and the precision/recall metrics were calculated using the feature space obtained by the VGG16 model. We use the UMAP-transformed space only for visualization.
@@
To evaluate whether the UMAP transformation effectively captures the distribution of data and the characteristics of the trajectories in the original space, we performed two analyses. First, we applied k-means clustering to the UMAP-transformed training data to confirm that UMAP projected the three categories into two-dimensional space while preserving their distinctions. The Adjusted Rand Score [31, 32], a measure of clustering accuracy, was approximately $0.976$. This high score indicates that UMAP successfully projected the clusters of the original categories into two dimensions while maintaining their differences. Next, we employed the Trustworthiness and Continuity metrics [33] to verify that the trajectories in the UMAP space effectively capture the high-dimensional dynamics. These metrics quantify the extent to which points in close proximity in high dimensions remain close in low dimensions, and vice versa. The metrics range from 0 to 1, with values closer to 1 indicating better preservation of the data structure. For the trajectories of our dynamical system, the Trustworthiness and Continuity metrics were approximately $0.942$ and $0.978$, respectively. These high values demonstrate that the UMAP representation effectively preserves the characteristics of the trajectories from the high-dimensional space.
@@
To quantitatively evaluate the quality and diversity of the generated image sequences, we calculated the precision and recall [19] (P/R) of the generated images from the MNIST-trained model using the method described in the subsection entitled “Precision and recall.” Here, quality refers to how similar the generated images are to the real dataset images, whereas diversity refers to how well the generated images cover a wide range of variations in the real dataset. The important parameters in this method are the number of data samples and the number of nearest neighbors. In this method, an equal number of real and generated data samples should be prepared. In our case, the real data consisted of all the test dataset images from categories $X$, $Y$, and $Z$, amounting to $3178$ samples. For the generated images, transformations were repeatedly applied starting from a test image $x_{0}$ selected from category $X$ as an initial value, resulting in a trajectory of 3178 points $\{x_{1},\ldots,x_{3178}\}$, and these samples included in this trajectory were used for evaluation. Previous research [19] proposed using $2\times 10^{4}$ samples each for real and generated data as a standard procedure. However, in this case, the number of test images does not reach this number, so direct comparison with other literature results is impossible. To ensure the validity of the evaluation, $3178$ samples from the same categories were extracted from the training dataset, and the P/R between the training and test datasets was calculated as a reference value. In the literature [19], the number of nearest neighbors $k$ is typically set to 3. However, because the number of data samples used in this study was different from that in the literature, it is difficult to draw conclusions from results with only a single value of $k$. Therefore, we varied $k$ from 1 to 10 and observed the resulting changes in P/R values.
@@
To investigate the dependence of the number of generated sequences on the evaluation of P/R values, we performed an additional analysis using a large number of artificially generated initial images and their trajectories (S4 Fig.). While the P/R values changed quantitatively due to the increased number of data points, the overall trends remained consistent. Specifically, we observed persistent high Precision/low Recall patterns and an initial decrease in Recall over time.
@@
S5 Fig. shows the histograms of the first five Lyapunov exponents calculated for each of the 980 trajectories. The histograms exhibit unimodal Gaussian-like distributions, suggesting that these trajectories converge to the same attractor without multi-stability. Therefore, it is reasonable to average these values to estimate the Lyapunov exponents.
Fig 8 shows all the spectrum values, with the inset showing an enlarged view of the first 15 Lyapunov exponents. The first seven Lyapunov exponents are definitely positive, and the largest Lyapunov exponent is estimated to be about 0.340. The presence of these positive Lyapunov exponents indicates that the generated trajectories exhibit sensitive dependence on initial conditions and are chaotic, suggesting that the chaotic attractor generates various images. Most Lyapunov exponents are negative except for the seven largest. The phenomenon of most exponents becoming negative is commonly observed in large-scale, high-dimensional dynamical systems where the dimension of the attractor is considerably smaller than the dimension of the system’s phase space [13, 34, 35].
@@
Further, S6 Fig. demonstrates the convergence of estimated Lyapunov exponents over time, obtained by calculating trajectories from 10 initial values for an extended period ($2\times 10^{6}$ steps). All ten trajectories converge to almost the same value, which matches the average value shown in Fig 8. This result provides an additional evidence for the stability of our Lyapunov exponent calculations in this model.
@@ Discussion
We conducted a visual investigation of the generated images to address whether the samples within the attractor possess any specific features that render them distinct from the samples outside the attractor. However, it was difficult to characterize the images within the attractor as either typical or atypical samples of their respective categories.
@@
This emergence of the large chaotic attractor may be related to our dynamical system’s design, which learned a cyclic path using only the generator $G$, without involving the other generator $F$. For comparison, consider a dynamical system constructed using the original CycleGAN [4], defined as $x_{n+1}=F(G(x_{n}))$. In this case, the composition $F\circ G$ is trained to approximate the identity mapping. Consequently, the trajectory is expected to converge to a small region, such as a fixed point, where no significant changes in the image are expected. In contrast, we constructed a system that circulates among three categories using only $G$, without being pulled back by $F$. We hypothesize that one of the primary reasons for the emergence of the large chaotic attractor is that this structure is not constrained by the requirement to approach the identity map. Further comparison of our model’s dynamics with those of CycleGAN, and a deeper understanding of the mechanisms underlying the emergence of the chaotic attractor, remain important topics for future research.
@@
To investigate the scope of our method with more challenging real-world examples, we believe it is necessary to extend the approach to handle transformations in latent space, similar to techniques used in StyleGAN [40] or latent diffusion models [41]. To link our method to associative memory processes and achieve more flexible image transformations, several advancements are needed. Potential avenues for future research include: a) Extending the method to allow variation of specific features within a category (e.g., the thickness or inclination of a character) through external conditional inputs. b) Developing the capability to store and navigate multiple category cycles. c) Adding functionality to dynamically change the association target based on conditional inputs.
@@ Acknowledgments
Precision and recall as functions of time step, calculated from trajectories from perturbed initial images. To investigate the dependence of the number of generated sequences on the evaluation of P/R values, we conducted an analysis using a large number of artificially generated initial images and their trajectories. We perturbed the initial images using the same method as in our maximum Lyapunov exponent analysis (Fig 9) and generated $15735$ trajectories ($3147\times 5$). The strength of the perturbation $\varepsilon$ was set to $0.1$. We then evaluated the Precision/Recall metrics using these trajectories and perturbed test images following the same procedure as in Fig 7.
S5 Fig.
@@
S6 Fig.
Convergence of estimated Lyapunov exponents over time.